4x+8y=2
8x-4y=4
ARE THESE LINES PERPENDICULAR, PARALLEL, OR NETHER?

Answer by Alan3354(69443)

4x+8y=2
8x-4y=4
ARE THESE LINES PERPENDICULAR, PARALLEL, OR NETHER?
---------------------------
Find the slopes of the lines. If the slopes are the same, they're parallel. If one slope is the negative inverse of the other, they're perpendicular. Otherwise, it's neither.
-------------
Find the slope by putting them in the slope-intercept form, y = mx + b.
In other words, solve for y.
4x+8y=2
8y = -4x + 2
y = (-1/2)x + 1/4
m1 = -1/2
-------------
8x-4y=4
-4y = -8x + 4
y = 2x - 1
m2 = +2
So they're perpendicular, since m1 = -1/m2